Grade 12Physical Chemistry

Four faradays of electricity were passwd thru AgNO3, CdSO4.AlCl3 and PbCl4 kept in four vessels using inert electrodes.The ratio of moles of Ag,Cd, Al and Pb deposited will be.

Shriya Mehrotra
11 Years agoGrade 12

4 Answers

Ramreddy
11 Years ago
This is according to Faradays laws of electrolysis i.e. W/E= constant
WAg / EAg = WCd / ECd = WAl / EAl =WPb / EPb = constant
EAg = MAg ;ECd = MCd /2 ;EAl = MAl /3;EPb = MPb /2
n Ag = 2 *nCd = 3* n Al = 2 *nPb
Supriyo saha
9 Years ago
Ag+=1,Cd2+=2,Al3+=3,Pb4+=4Four faraday will depisite 4*1=4moles of Ag, (4÷2)moles of Cd, (4÷3)moles of Al, and (4÷4)=1moles of PbAg:Cd:Al:Pb=4:(4÷2):(4÷3):1 =4:2:(4÷3):1 =12:6:4:3
Lucio Demelo
9 Years ago
What is the first answer!! Makes no sense waste of people`s the reading it please do check what people are posting and then allow that answer to be viewed
ankit singh
5 Years ago
n Ag = 2 *nCd = 3* n Al =/2Pb = MPb/3;EAl = MAl/2 ;ECd = MCd ;EAg= MAg = constant EPb / EPb =WAl / EAl = WCd / ECd = WAg / EAg This is according to Faradays laws of electrolysis i.e. W/E= constant W
Ag+=1,Cd2+=2,Al3+=3,Pb4+=4Four faraday will depisite 4*1=4moles of Ag, (4÷2)moles of Cd, (4÷3)moles of Al, and (4÷4)=1moles of PbAg:Cd:Al:Pb=4:(4÷2):(4÷3):1 =4:2:(4÷3):1 =12:6:4:3